# {cutout}02.753cm0cm1 empty ViC-Bench     : Benchmarking Visual-Interleaved Chain-of-Thought Capability in MLLMs with Free-Style Intermediate State Representations

### v2

@@ Abstract
Visual-Interleaved Chain-of-Thought (VI-CoT) enables MLLMs to continually update their understanding and decisions based on step-wise intermediate visual states (IVS), much like a human would, which demonstrates impressive success in various tasks, thereby leading to emerged advancements in related benchmarks. Despite promising progress, current benchmarks provide models with relatively fixed IVS, rather than free-style IVS, whch might forcibly distort the original thinking trajectories, failing to evaluate their intrinsic reasoning capabilities. More importantly, existing benchmarks neglect to systematically explore the impact factors that IVS would impart to untamed reasoning performance. To tackle above gaps, we introduce a specialized benchmark termed ViC-Bench, consisting of four representive tasks: maze navigation, jigsaw puzzle, embodied long-horizon planning, and complex counting, where each task has dedicated free-style IVS generation pipeline supporting function calls. To systematically examine VI-CoT capability, we propose a thorough evaluation suite incorporating a progressive three-stage strategy with targeted new metrics. Besides, we establish Incremental Prompting Information Injection (IPII) strategy to ablatively explore the prompting factors for VI-CoT. We extensively conduct evaluations for 18 advanced MLLMs, revealing key insights into their VI-CoT capability. Our proposed benchmark is publicly open at [Uncaptioned image] Huggingface.
11footnotetext: Work done during research internship at Meituan-M17.22footnotetext: Equal Contribution.33footnotetext: Project Lead.
1 Introduction
@@
AI field is currently evolving from LLMs [45, 9, 69] to MLLMs [8, 26, 74, 72], which integrate various modalities into the language decoders [73]. Achieving human-level multi-modal intelligence requires transcending basic perceptual capabilities to attain sophisticated reasoning. Drawing inspirations from the remarkable success of Chain-of-Thought (CoT) in LLMs [11, 38, 24], the integration of CoT into visual-language contexts has catalyzed transformative progress, giving rise to visual CoT [33, 73].
@@
The initial visual CoT involves vision signals only as input, whereas the entire rationales are composed of language, in which various methods and related benchmarks make rapid advancements [10, 6, 77]. However, this paradigm overlooks the explicit visual representation updates and continuous understanding of visual feedbacks, misaligning with the human cognitive process of using visual thoughts for concrete reasoning and textual thoughts for abstract reasoning. To this end, Visual-Interleaved Chain-of-Thought (VI-CoT), which incorporates step-wise intermediate visual states (IVS) based on visual inputs, has made rapid progress. According to the source of IVS, current VI-CoT methods are primarily divided into two types. The first type involves autonomously generating IVS based on its internalized understanding [76, 33, 84]. However, this approach currently struggles due to the limited generative capabilities of MLLMs [76, 84]. The second type involves providing IVS through external knowledge retrieval, utilizing expert tools or human in the agent-form, which shows impressive results in broad tasks [81, 13, 34]. Meanwhile, to evaluate the developments of recent VI-CoT methods, various benchmarks have emerged [13, 82]. Despite promising advancements, few of them provides free-style IVS representations to MLLMs, as illustrated in Tab. 1. CoMT [13] primarily provides fixed IVS, which might forcibly distort the original planning trajectories. While MageBench [82] offers the dynamic IVS but imposes the attribute constraints of action-observation memory. More importantly, existing benchmarks [60, 28, 13, 6] neglect to systematically assess the impact factors that IVS would impart to untamed reasoning performance in MLLMs. (i.e., Positive, Negative, or Null). As a result, a natural question arise: Could MLLMs leverage VI-CoT, which closely aligns with human cognitive behavior, to inherently achieve better reasoning performance?
@@
Figure 1: Performance overview of advanced MLLMs on four tasks in average ACC across stages.
@@
To tackle above gaps, we introduce the specialized ViC-Bench for evaluating VI-CoT, selecting four representative tasks (i.e., maze navigation, jigsaw puzzle, embodied long-horizon planning, and complex counting), which require models to dynamically interact with visual contexts and continuously update their understanding and decision-making based on step-wise IVS. We first propose dedicated data construction pipeline, resulting in 250 unique images for each task. We then introduce free-style IVS generation workflows supporting function calls to sufficiently support the investigation on VI-CoT performance. Based on the constructed data, we propose a novel evaluation suite incorporating progressive three-stage evaluation strategy with targeted new metrics. Specifically, Stage 1 involves multiple-choice QA, while Stage 2 focuses on open-ended QA, both utilizing visual signals solely at the input to establish solid foundations for Stage 3. Building on this, Stage 3 features open-ended QA with free-form IVS, enriching from the hierarchical references for VI-CoT evaluation through our progressive design. Subsequently, we define new Recall and ThinkGain metrics by black-boxing the retrospection process, along with a Legality metric to tackle the clear rule boundaries, thereby establishing a comprehensive suite for each stage. To ablatively explore the prompting factors affecting VI-CoT capability, we further design the Incremental Prompting Information Injection (IPII) strategy across three stages, utilizing varying global-aware prompting levels. We totally examine 18 SOTA MLLMs on our ViC-Bench, offering extensive quantitative and qualitative results, as illustrated in Fig. 1. We uncover significant performance gaps between open-source and proprietary MLLMs in both quantitative analyses and prompting studies. Moreover, most MLLMs show substantial disparities compared to human-level proficiency. For example, powerful Gemini-2.5-pro [63] performs inadequately on jigsaw puzzle, only achieving 6.80% ACC in Stage 2.
@@
In addition to providing insights into VI-CoT capabilities of MLLMs, ViC-Bench can also establish foundations for the developments of native MLLMs [59, 61], multi-modal agents [79, 56], embodied AI [34, 35], and autonomous driving [15]. The main contributions of this paper are three-fold:
@@
    We curate ViC-Bench, including four representative tasks, each with dedicated construction and free-style IVS generation pipelines. Moreover, we engage human-machine collaborations in both construction and evaluation to deliver a high-quality benchmark for VI-CoT reasoning.
@@
    We propose a thorough evaluation suite that includes a progressive three-stage evaluation strategy with newly targeted metrics to meticulously examine inherent VI-CoT performance using free-style IVS. We further introduce the IPII strategy to ablatively explore the prompting factors for VI-CoT.
@@
    Extensive experiments and analyses are conducted on 18 proprietary and open-source MLLMs. We summarize several key observations and insights, hoping to inspire advancements of future works.
Table 1: The statistics comparisons of ViC-Bench and related representative benchmarks. #No.: Unique sample number. Free: Free-style IVS supporting function calls.
@@
M3CoT [10] (Test Part) ACL’24 Web 2,358 3 $\times$ $\times$
LEGO-Puzzles [60] arXiv’25 Synthesized 1,100 3 $\times$ $\times$
MME-CoT [28] ICML’25 Web 808 6 $\times$ $\times$
CoMT [13] AAAI’25 Web 3,853 4 $\times$
MageBench [82] arXiv’24 Synthesized & Web 483 3
VERIFY [6] arXiv’25 Web 600 1 $\times$ $\times$ $\times$
@@
2 Related Work
@@
Multi-modal Large Language Models. The integration of multi-modal information with LLMs [45, 9] leads to the emergence of MLLMs [8, 26, 74, 66], exhibiting impressive performance in various multi-modal understanding and generation tasks. MLLMs can be broadly categorized into two types: pipeline-based and native paradigms. The pipeline-based MLLMs can be classified into three types based on the multi-modal integration strategies: (1) Feature mapping with MLPs, such as PaLM-E [19], LLaVA [39], and CogVLM [71]; (2) Query-based cross-attention components (e.g., InstructBLIP [16], Mini-GPT4 [86], Qwen-VL series [4, 70, 5]); (3) Cross-attention layers within LLMs, such as Flamingo [1] and IDEFICS [30]. Meanwhile, MLLMs integrated with generation capabilities through coupled components have also been largely promoted [51, 59, 85]. As for native MLLMs, they primarily achieve unified understanding and generation through auto-regressive manners with elaborate tokenizers [75, 40, 61]. Most recently, the release of OpenAI o3/o4 [49] and DeepSeek-R1 [24] sparks a wave of interests in reasoning enhancements, highlighting the effectiveness of CoT [35, 80]. Inspired by this, researchers advance reasoning capabilities of MLLMs employing visual CoT mechanisms [43, 55, 41].
@@
Visual-Interleaved Chain-of-Thought. VI-CoT involves the engagement of step-wise IVS through the entire reasoning process, achieving impressive performance across various downstream scenarios. However, due to limited generative capabilities, MLLMs struggle to generate the native IVS, which are essential for the in-context knowledge retrieval. As a result, current methods primarily rely on external knowledge retrieval to develop IVS, such as expert tools or human in the agent-form [73, 38]. CMMCoT [81] utilizes the visual region tokens as supervisory signals to perform interleaved reasoning. Zhang et al. [83] extend o1-style reasoning to interactive embodied search. Gao et al. [23] propose the attention-driven selection method to realize interleaved CoT. VoT [21] breaks down complex task into sub-problems, and address them from low to high using scene graphs. MVoT [33] enables visual thinking by generating visual visualizations of reasoning traces. Hu et al. [25] provide MLLMs with sketchpad and expert tools to conduct interleaved CoT. Meanwhile, related benchmarks have emerged to extensively evaluate various VI-CoT methods. CoMT [13] constructs four types of visual operations, requiring multi-modal reasoning outputs. Zhang et al. [82] propose MageBench for evaluating the MLLMs’s capabilities of being an agent. [83] cultivates 809 test cases across 12 scenarios for hierarchical long-horizon tasks. Despite great advancements, few of these benchmarks provide the free-form IVS representations and systematically evaluate the impacts that IVS might have on the untamed reasoning performance. To bridge these gaps, we carefully construct four representative VI-CoT tasks with the free-form IVS representations and further propose a comprehensive evaluation suite incorporating three stages.
3 ViC-Bench Construction
3.1 Overview
@@
We introduce ViC-Bench, consisting of four representative VI-CoT tasks: Maze Navigation (Sec. 3.2), Jigsaw Puzzle (Sec. 3.3), Embodied Long-Horizon Planning (Sec. 3.4), and Complex Counting (Sec. 3.5). The overall construction workflow can be mainly divided into Raw Data & Pre-processing, Three-Stage Construction, IVS Generation, and Human Recheck, as displayed in Fig. 2.
3.2 Maze Navigation
@@
Raw Data & Pre-processing. As shown in Fig. 2 (a), we utilize Maze [27] library coupled with DFS method to render 4 $\times$ 4 mazes in multiple batches. After generation, we first screen out those with navigation lengths ranging from 5-8, then aggregate mazes with the same starting point. Finally, mazes with duplicate shortest path after global-aware aggregation are removed to ensure uniqueness.
@@
Stage 1. For the processed mazes, we randomly select an original maze $\mathbf{M_{1}}$ in the non-overlapping paradigm, then select three other original mazes $\mathbf{M_{2}}$, $\mathbf{M_{3}}$, and $\mathbf{M_{4}}$. Subsequently, we draw the endpoints of $\mathbf{M_{2}}$-$\mathbf{M_{4}}$ onto $\mathbf{M_{1}}$ to establish three incorrect options, facilitating MLLMs to select the correct endpoint based on the starting point and the given navigation path.
@@ 3.3 Jigsaw Puzzle
Raw Data & Pre-processing. To eliminate the risk of data leakage from the familiar datasets [17, 36], we construct an image source pool using elaborate prompts with DALLE-3 [46], FLUX.1-schnell [29], Kwai-Kolors [65], Stable-Diffusion-3 Medium [52], WanX [2], and Midjourney-V6.1 [44], which can sufficiently ensure the image diversity and uniqueness. Following [31, 37], we employ the manual white-box approach to filter the generated images, primarily considering three discriminative metrics: T2I consistency, reasonableness, as well as realism, i.e.,
$S_{overall}=0.6\times S_{consistency}+0.2\times S_{reasonableness}+0.2\times S_{realism},$ (1)
@@
where $S\in[1,5]\cap\mathbb{Z}$. Based on the descending order of overall scores, we finally select 250 images.
@@
Stage 1. We first adjust the selected images to 224 $\times$ 224 resolutions and divide them into 4 $\times$ 4 patches. We then targetedly extract six patches using proposed weighted dispersed sampling strategy. Afterwards, the selected patches are removed from the original image to generate masked image $\mathbf{I}_{m}$. Subsequently, the chosen patches are randomly numbered and concatenated with $\mathbf{I}_{m}$ along the vertical direction to output the overall input image, as illustrated in Fig. 2 (b). As for options, four patches are first randomly selected and correctly positioned to establish the correct option. Three incorrect options are then generated based on the puzzle states with either (3 ✓& 1 $\times$) or (2 ✓& 2 $\times$) patch placements, leading to the plausible yet incorrect paradigm, which challenge the discriminative capabilities of MLLMs in region-aware semantic understanding and spatial-aware reasoning.
@@
Stages 2 & 3. The overall input images of last two evaluation stages almost keep the same with Stage 1, but additionally label the six vacant regions in $\mathbf{I}_{m}$. Moreover, the arrangement of six patches in the overall input image underneath remains consistent with Stage 1. The task instructions here require models to directly output the correct mapping between patches and vacant regions.
@@
IVS Generation. Based on the simulated functions and planning trajectories of MLLMs, we conduct multi-step agent-form evaluations with free-style IVS. Besides, we establish the rule boundaries such that if a patch or vacant region is repeatedly utilized, causing conflicts, we will deem the current action invalid and provide appropriate guidance. We also set the maximum attempts to 30.
@@
Human Recheck. Following maze navigation, we also employ human experts to recheck the generated puzzles across stages and perform quality inspections on VI-CoT evaluations in Stage 3.
@@
Figure 2: The illustrations of the overall construction pipelines for four downstream tasks.
3.4 Embodied Long-Horizon Planning
Raw Data & Pre-processing. We establish this task based on the benchmark from Embodied-reasoner [83]. We first consider the length of action sequences for manual selection (i.e., longer than 20 steps are removed) and further ensure the uniqueness of each sequence. Semantic aggregations are then performed to promote sequence separation. In the end, we manually select 250 samples.

### v3

# ViC-Bench: Benchmarking Visual-Interleaved Chain-of-Thought Capability in MLLMs with Free-Style Intermediate State Representations

@@ Abstract
Visual-Interleaved Chain-of-Thought (VI-CoT) enables Multi-modal Large Language Models (MLLMs) to continually update their understanding and decision space based on step-wise intermediate visual states (IVS), much like a human would, which has demonstrated impressive success in various tasks, thereby leading to emerged advancements in related downstream benchmarks. Despite promising progress, current benchmarks provide models with relatively fixed IVS, rather than free-style IVS, whch might forcibly distort the original thinking trajectories, failing to evaluate their intrinsic reasoning capabilities. More importantly, existing benchmarks neglect to systematically explore the impact factors that IVS would impart to the untamed reasoning performance. To tackle above gaps, we introduce a specialized benchmark termed ViC-Bench, consisting of four representive tasks, i.e., maze navigation, jigsaw puzzle, embodied long-horizon planning, as well as complex counting, where each task has dedicated free-style IVS generation pipeline supporting adaptive function calls. To systematically examine VI-CoT capability, we propose a thorough evaluation suite incorporating a progressive three-stage strategy with targeted new metrics. Besides, we establish Incremental Prompting Information Injection strategy to ablatively explore the prompting factors for VI-CoT. We extensively conduct evaluations for 18 advanced MLLMs, revealing key insights into their VI-CoT capability. The introduced ViC-Bench has been made publicly available at Huggingface.
I Introduction
@@
Multi-modal AI field is currently evolving from LLMs [LLM_review-1, LLM_review-2, Llama3_touvron2023llama] to MLLMs [MLLM_review-1, TIP-VLMs-2, TIP-VLMs], which integrate various modalities into the backend language decoders [Multimodal-chain-of-thought]. Achieving human-level multi-modal intelligence requires transcending basic perceptual capabilities to attain sophisticated reasoning. Drawing inspirations from the remarkable success of Chain-of-Thought (CoT) in LLMs [Towards-reasoning-era, Mind-with-eyes, Deepseek-r1], the integration of CoT into visual-language contexts has catalyzed transformative progress, giving rise to visual CoT [li2025imagine-CoT, Multimodal-chain-of-thought].
@@
The initial visual CoT involves vision signals only as input, whereas the entire rationales are composed of language, in which various methods and related benchmarks make rapid advancements [Multi-Evaluation, VERIFY, xu2024LLaVA-CoT]. However, this paradigm overlooks the explicit visual representation updates and continuous understanding of visual feedbacks, misaligning with the human cognitive process of using visual thoughts for concrete reasoning and textual thoughts for abstract reasoning. To this end, Visual-Interleaved Chain-of-Thought (VI-CoT), which incorporates step-wise intermediate visual states (IVS) based on visual inputs, has made rapid progress. According to the source of IVS, current VI-CoT methods are primarily divided into two types. The first type involves autonomously generating IVS based on its internalized understanding [li2025imagine-CoT, RISEBench]. However, this approach currently struggles due to the limited generative capabilities of MLLMs [Multi-step, RISEBench]. The second type involves providing IVS through external knowledge retrieval, utilizing expert tools or human in the agent-form, which shows impressive results in various tasks [CMMCoT, Comt_AAAI, Embodied-agent-interface]. Meanwhile, to evaluate the developments of recent VI-CoT methods, various benchmarks have emerged [Comt_AAAI, magebench]. Despite promising advancements, few of them provides free-style IVS representations to MLLMs, as illustrated in Tab. I below. CoMT [Comt_AAAI] primarily provides fixed IVS, which could forcibly distort the original planning trajectories of models. While MageBench [magebench] offers the dynamic IVS but imposes the attribute constraints of action-observation memory. More importantly, existing benchmarks [LEGO-Puzzles, MME-CoT, Comt_AAAI, VERIFY] neglect to systematically assess the impact factors that IVS would impart to the untamed reasoning performance in MLLMs. (i.e., Positive, Negative, or Null). As a result, a natural question arises: Could MLLMs leverage VI-CoT, which closely aligns with human cognitive behavior, to inherently achieve better reasoning performance?
@@
Figure 1: Performance overview of advanced MLLMs on four tasks in terms of average ACC across three evaluation stages.
@@
To tackle above gaps, we introduce the specialized ViC-Bench for evaluating VI-CoT, selecting four representative tasks (i.e., maze navigation, jigsaw puzzle, embodied long-horizon planning, and complex counting), which require models to dynamically interact with visual contexts and continuously update their understanding and decision-making based on step-wise IVS. We first propose dedicated data construction pipelines, resulting in 250 unique images for each task. We then introduce free-style IVS generation workflows supporting function calls to sufficiently support the investigation on VI-CoT performance. Based on the constructed data, we propose a novel evaluation suite incorporating progressive three-stage evaluation strategy with targeted new metrics. Specifically, Stage 1 involves multiple-choice QA, while Stage 2 focuses on open-ended QA, both utilizing visual signals solely at the input to establish solid foundations for Stage 3. Building on this, Stage 3 features open-ended QA with free-form IVS, enriching from the hierarchical references for VI-CoT evaluation through our progressive design. Subsequently, we define new Recall and ThinkGain metrics by black-boxing the retrospection process, along with a Legality metric to tackle the clear rule boundaries, thereby establishing a comprehensive suite for each stage. To ablatively explore the prompting factors affecting VI-CoT capability, we further design the Incremental Prompting Information Injection (IPII) strategy across three stages, utilizing varying global-aware prompting levels. We totally examine 18 advanced MLLMs on introduced ViC-Bench, providing extensive quantitative and qualitative results. The performance overview is illustrated as Fig. 1 above. We uncover significant performance gaps between open-source and proprietary MLLMs in both quantitative analyses and prompting studies. Moreover, most MLLMs show substantial disparities compared to human-level proficiency.
@@
In addition to providing insights into VI-CoT capabilities of MLLMs, ViC-Bench can also establish foundations for the progress of unified MLLMs [Emu3d5, bagel], multi-modal agents [multi-modal-agent-1, multi-modal-agent-2], embodied AI [Embodied-agent-interface, system2-review1], and autonomous driving [autonomous-driving]. In summary, the main contributions of this paper are three-fold:
@@
    We curate ViC-Bench, including four representative tasks, each with dedicated construction and free-style IVS generation pipelines. Moreover, we engage human-machine collaborations in both construction and evaluation to establish a high-quality benchmark for VI-CoT reasoning.
@@
    We propose a thorough evaluation suite that includes a progressive three-stage evaluation strategy with newly targeted metrics to meticulously examine the inherent VI-CoT performance using free-style IVS. We further introduce the Incremental Prompting Information Injection (IPII) strategy to ablatively explore the prompting factors for VI-CoT.
@@
    Extensive experiments and analyses are performed on 18 proprietary and open-source MLLMs. We summarize several key observations and insights, hoping to inspire advancements of future research.
TABLE I: The statistics comparisons of ViC-Bench and related representative benchmarks. #No.: Unique sample number. Free-Style: Free-style IVS supporting function calls.
@@
M3CoT [m3cot] (Test Part) ACL’24 Web 2,358 3 $\times$ $\times$
LEGO-Puzzles [LEGO-Puzzles] arXiv’25 Synthesized 1,100 3 $\times$ $\times$
MME-CoT [MME-CoT] ICML’25 Web 808 6 $\times$ $\times$
CoMT [Comt_AAAI] AAAI’25 Web 3,853 4 $\times$
MageBench [magebench] arXiv’24 Synthesized & Web 483 3
VERIFY [VERIFY] arXiv’25 Web 600 1 $\times$ $\times$ $\times$
@@
II Related Work
II-A Multi-modal Large Language Models
@@
The integration of multi-modal information with LLMs [LLM_review-1, LLM_review-2] leads to the emergence of MLLMs [MLLM_review-1, TIP-VLMs-2, TIP-VLMs], exhibiting impressive performance in various multi-modal understanding and generation tasks. MLLMs can be broadly categorized into two types: pipeline-based and native paradigms. The pipeline-based MLLMs can be generally classified into three types based on the multi-modal integration strategies: (1) Feature mapping with MLPs, such as PaLM-E [30_driess2023palm], LLaVA [29_liu2024visual], and CogVLM [32_wang2023cogvlm]; (2) Query-based cross-attention components (e.g., InstructBLIP [4_dai2023instructblip], Mini-GPT4 [26_zhu2023minigpt], and Qwen-VL series [28_bai2023qwen, Qwen2-vl, Qwen2d5-vl]); (3) Cross-attention layers within LLMs, such as Flamingo [35_alayrac2022flamingo] and IDEFICS [37_laurenccon2024obelics]. Meanwhile, MLLMs integrated with generation capabilities through coupled components have also been largely promoted [44_peng2023kosmos, Emu, Minigpt-5]. As for native MLLMs, they primarily achieve unified understanding and generation through auto-regressive manners with elaborate tokenizers [Liquid, LWM2024, Chameleon]. Most recently, the release of OpenAI o3/o4 [openai2025o4mini] and DeepSeek-R1 [Deepseek-r1] sparks a wave of interests in reasoning enhancements, highlighting the effectiveness of CoT [system2-review1, zeng2024mr-Mr-ben]. Inspired by this, researchers have sought to advance the reasoning capabilities of MLLMs by employing visual CoT mechanisms. [MM-Eureka, Vlm-r1, liu2024enhancing-CoT-ZJU].
II-B Visual-Interleaved Chain-of-Thought
@@
VI-CoT involves the engagement of step-wise IVS through the entire reasoning process, achieving impressive performance across various downstream scenarios. However, due to the limited visual generative capabilities, MLLMs struggle to generate the native IVS, which are essential for the in-context knowledge retrieval. As a result, current methods primarily rely on external knowledge retrieval to develop IVS, such as expert tools or human in the agent-form [Multimodal-chain-of-thought, Mind-with-eyes]. CMMCoT [CMMCoT] utilizes the visual region tokens as supervisory signals to perform interleaved reasoning. Zhang et al. [Embodied-reasoner] extend o1-style reasoning to interactive embodied search. Gao et al. [gao2024interleaved-CoT] propose the attention-driven selection method to realize interleaved CoT. VoT [Video-of-thought] breaks down complex task into sub-problems, and address them from low to high employing scene graphs. MVoT [li2025imagine-CoT] enables visual thinking by generating visual visualizations of reasoning trajectories. Hu et al. [Visual-sketchpad] provide MLLMs with sketchpad and expert tools to conduct interleaved CoT. Meanwhile, related benchmarks have emerged to extensively evaluate various VI-CoT methods. CoMT [Comt_AAAI] constructs four types of visual operations, requiring multi-modal reasoning outputs. Zhang et al. [magebench] propose MageBench for evaluating the MLLMs’s capabilities of being an agent. [Embodied-reasoner] totally cultivates 809 test cases across 12 scenarios for hierarchical embodied long-horizon tasks. Despite great advancements, few of these benchmarks provide the free-form IVS representations and systematically evaluate the influence that IVS can make on the untamed reasoning performance. To bridge these gaps, we carefully construct four representative VI-CoT tasks with the free-form IVS representations and further propose a comprehensive evaluation suite incorporating three progressive stages.
III Benchmark Construction
III-A Overview
@@
We introduce ViC-Bench, comprising four representative VI-CoT tasks, i.e., Maze Navigation (Sec. III-B), Jigsaw Puzzle (Sec. III-C), Embodied Long-Horizon Planning (Sec. III-D), and Complex Counting (Sec. III-E). The overall construction workflow can be mainly divided into Raw Data & Pre-processing, Three-Stage Construction, IVS Generation, and Human Recheck, as illustrated in Fig. 2.
III-B Maze Navigation
@@
Raw Data & Pre-processing. As shown in Fig. 2 (a), we utilize Maze [maze-dataset] library coupled with DFS method to render 4 $\times$ 4 mazes in multiple batches. After generation, we first screen out those with navigation lengths ranging from 5-8, then aggregate mazes with the same starting point. Finally, mazes with duplicate shortest path after global-aware aggregation are removed to ensure uniqueness.
@@
Stage 1. For the processed mazes, we randomly select an original maze $\mathbf{M_{1}}$ in the non-overlapping paradigm, then select three other distinct mazes $\mathbf{M_{2}}$, $\mathbf{M_{3}}$, and $\mathbf{M_{4}}$. Subsequently, we draw the endpoints of $\mathbf{M_{2}}$-$\mathbf{M_{4}}$ onto $\mathbf{M_{1}}$ to establish three incorrect options serving as visual distractors, facilitating MLLMs to select the correct endpoint based on the starting point and the given navigation path.
@@ III-C Jigsaw Puzzle
Raw Data & Pre-processing. To eliminate the risk of data leakage from the familiar datasets [56_deng2009imagenet, 57_lin2014microsoft], we construct an image source pool using elaborate prompts with DALLE-3 [DALLE-3], FLUX.1-schnell [flux2024], Kwai-Kolors [kolors], Stable-Diffusion-3 Medium [stablediffusion], WanX [WanX], and Midjourney-V6.1 [Midjourney], which can sufficiently ensure the image diversity and uniqueness. Following [t2i-eva-1, t2i-eva-2], we employ the manual white-box approach to filter the generated images, primarily considering three discriminative metrics, i.e., T2I consistency, reasonableness, and, which can be formulated as:
$S_{overall}=0.6\times S_{cy}+0.2\times S_{rs}+0.2\times S_{rm},$ (1)
@@
where $S_{cy}$, $S_{rs}$, and $S_{rm}$ denote the scores of consistency, reasonableness, as well as realism, respectively. Moreover, $S\in[1,5]\cap\mathbb{Z}$. Based on the descending order of overall scores, we finally select 250 unique images.
@@
Stage 1. We first adjust the selected images to 224 $\times$ 224 resolutions and divide them into 4 $\times$ 4 patches. We then targetedly extract six patches using proposed weighted dispersed sampling strategy (as described in Sec. IV of supplementary material). Afterwards, the selected patches are removed from the original image to generate masked image $\mathbf{I}_{m}$. Subsequently, the chosen patches are randomly numbered and concatenated with $\mathbf{I}_{m}$ along the vertical direction to output the overall input image, as illustrated in Fig. 2 (b). As for options, four patches are first randomly selected and correctly positioned to establish the correct option. Three incorrect options are then generated based on the puzzle states with either (3 ✓& 1 $\times$) or (2 ✓& 2 $\times$) patch placements, leading to the plausible yet incorrect paradigm, which challenge the discriminative capabilities of MLLMs in region-aware semantic understanding and cross-level spatial reasoning.
@@
Stages 2 & 3. The overall input images of the last two evaluation stages almost keep the same with Stage 1, but additionally label the six vacant regions in $\mathbf{I}_{m}$. Moreover, the arrangement of six patches in the overall input image underneath remains consistent with Stage 1. The task instructions here require models to directly output the correct mapping between patches and vacant regions.
@@
IVS Generation. Based on the simulated functions and planning trajectories of MLLMs, we conduct multi-step agent-form evaluations with free-style IVS. Besides, we establish the rule boundaries such that if a patch or vacant region is repeatedly utilized, causing conflicts, we will deem the current action invalid and provide appropriate guidance. To keep consistent with maze navigation, we set the maximum attempts to 30.
@@
Human Recheck. Following maze navigation, we also employ human experts to recheck the generated puzzles across all stages and perform thorough quality inspections on VI-CoT evaluations in Stage 3.
@@
Figure 2: The illustrations of the overall construction pipelines for four representative VI-CoT tasks in proposed ViC-Bench.
III-D Embodied Long-Horizon Planning
Raw Data & Pre-processing. We construct the dataset for this task based on the AI2-THOR simulation environment [ai2thor]. To ensure the task difficulty is appropriately calibrated for long-horizon planning, we filter action sequences based on trajectory length, retaining only those with a step count in the range of $3$ to $20$. Furthermore, to mitigate the issue of unreachable target locations caused by navigation errors within the simulator, we rigorously exclude search-oriented tasks. Specifically, tasks involving “PickUp”, “Put”, and other long-range navigation objectives are systematically removed to guarantee the validity and plausibility of the ground truth paths. In the end, we manually curate a diverse and high-quality subset of 250 samples for final evaluation.